# BVI-AOM: A New Training Dataset for Deep Video Compression Optimization

### v2

@@ Abstract
Deep learning is now playing an important role in enhancing the performance of conventional hybrid video codecs. These learning-based methods typically require diverse and representative training material for optimization in order to achieve model generalization and optimal coding performance. However, existing datasets either offer limited content variability or come with restricted licensing terms constraining their use to research purposes only. To address these issues, we propose a new training dataset, named BVI-AOM, which contains 956 uncompressed sequences at various resolutions from 270p to 2160p, covering a wide range of content and texture types. The dataset comes with more flexible licensing terms and offers competitive performance when used as a training set for optimizing deep video coding tools. The experimental results demonstrate that when used as a training set to optimize two popular network architectures for two different coding tools, the proposed dataset leads to additional bitrate savings of up to 0.29 and 2.98 percentage points in terms of PSNR-Y and VMAF, respectively, compared to an existing training dataset, BVI-DVC, which has been widely used for deep video coding. The BVI-AOM dataset is available for download under this link: (TBD).
@@ I Introduction
More recently, both MPEG and AOM have initialized new working models, in order to achieve further coding gains over the latest standards, H.266/VVC and AV1. In these models, in addition to sophisticated modifications to conventional coding modules, there are active investigations on the use of various deep learning techniques to obtain more evident improvement [7, 8, 9, 10, 11]. In parallel, deep neural networks have also been employed to build new coding frameworks [12, 13, 14] which enable end-to-end optimization. Although these neural codecs typically require additional graphics processing resources and are associated with high computational complexity, they do show great potential to compete with standard video coding algorithms.
@@ V Conclusion
In this paper, we present a new training dataset, BVI-AOM, for training deep video coding methods. It contains 956 uncompressed sequences with various spatial and temporal resolutions and offers a relatively wide coverage of low-level video features and texture types. When used for training various deep video coding tools, the results show that the BVI-AOM dataset offers consistent performance gains when compared to the commonly used dataset, BVI-DVC, with up to 2.98p.p. additional BD-rate saving. The proposed dataset comes with flexible licensing terms permitting its use for academic research and video standards development purposes. The dataset will be made publicly available once this paper is accepted.

### v3

@@ Abstract
Deep learning is now playing an important role in enhancing the performance of conventional hybrid video codecs. These learning-based methods typically require diverse and representative training material for optimization in order to achieve model generalization and optimal coding performance. However, existing datasets either offer limited content variability or come with restricted licensing terms constraining their use to research purposes only. To address these issues, we propose a new training dataset, named BVI-AOM, which contains 956 uncompressed sequences at various resolutions from 270p to 2160p, covering a wide range of content and texture types. The dataset comes with more flexible licensing terms and offers competitive performance when used as a training set for optimizing deep video coding tools. The experimental results demonstrate that when used as a training set to optimize two popular network architectures for two different coding tools, the proposed dataset leads to additional bitrate savings of up to 0.29 and 2.98 percentage points in terms of PSNR-Y and VMAF, respectively, compared to an existing training dataset, BVI-DVC, which has been widely used for deep video coding. The BVI-AOM dataset is available at https://github.com/fan-aaron-zhang/bvi-aom.
@@ I Introduction
More recently, both MPEG and AOM have initialized new working models, in order to achieve further coding gains over the latest standards, H.266/VVC and AV1. In these models, in addition to sophisticated modifications to conventional coding modules, there are active investigations on the use of various deep learning techniques to obtain more evident improvement [7, 8, 9, 10, 11, 12]. In parallel, deep neural networks have also been employed to build new coding frameworks [13, 14, 15] which enable end-to-end optimization. Although these neural codecs typically require additional graphics processing resources and are associated with high computational complexity, they do show great potential to compete with standard video coding algorithms.
@@ V Conclusion
In this paper, we present a new training dataset, BVI-AOM, for training deep video coding methods. It contains 956 uncompressed sequences with various spatial and temporal resolutions and offers a relatively wide coverage of low-level video features and texture types. When used for training various deep video coding tools, the results show that the BVI-AOM dataset offers consistent performance gains when compared to the commonly used dataset, BVI-DVC, with up to 2.98p.p. additional BD-rate saving. The proposed dataset comes with flexible licensing terms permitting its use for academic research and video standards development purposes. The dataset is publicly available under this link: https://github.com/fan-aaron-zhang/bvi-aom.